OpenStudy (anonymous):

Prove this identity: csc2 ϴ - cot2 ϴ = sec2 ϴ - tan2 ϴ PLEASE HELP ME! :D

OpenStudy (anonymous):

\[\frac{ 1 }{ \sin^2\theta }-\frac{ \cos^2\theta }{ \sin^2\theta }=\frac{ 1 }{ \cos^2\theta }-\frac{ \sin^2\theta }{ \cos^2\theta }\] \[LHS=\frac{ 1-\cos^2\theta}{\sin^2\theta }=\frac{ \sin^2\theta }{\sin^2\theta }=1\] \[RHS=\frac{ 1-\sin^2\theta }{ \cos^2\theta }=\frac{ \cos^2\theta }{ \cos^2\theta }=1\] RHS=LHS

OpenStudy (anonymous):

thanks but i already know the answer po. hehe . :D

OpenStudy (anonymous):

yw